# LocalSearchBench: Benchmarking Agentic Search in Real-World Local Life Services

### v1

@@ 5.3. Benchmark Results (RQ2)
The choice of underlying LRM significantly influences performance across all metrics. Among all models, DeepSeek-V3.1 achieves the highest Correctness (34.34%), followed by GLM-4.5 (33.78%) and LongChat-Large-32K (33.19%). Qwen-Plus-Latest leads in Completeness (80.94%), while Gemini-2.5-Pro excels in Faithfulness (78.26%). Closed-source models outperform open-source ones in completeness (79.08% vs. 75.57%) and efficiency (2.25 vs. 2.61 tool calls), though differences in Correctness are more modest. Despite these variations, overall performance remains limited, as even the best model achieves only 34.34% Correctness, underscoring the substantial challenge of multi-hop local search.
Table 5. Performance comparison of different models on LocalSearchBench (Max Round N = 5).
Model Config Evaluation Metrics
Avg. tool calls Avg. rounds Correctness (%) Completeness (%) Fluency (%) Faithfulness (%) Safety (%)
Closed-source LRMs
[Uncaptioned image] GPT-4.1 w/o web search 1.05 2.06 18.73 70.64 72.37 75.10 83.71
LocalPlayground 1.72 ($\uparrow$0.67) 2.70 ($\uparrow$0.64) 26.76 ($\uparrow$8.03) 75.42 ($\uparrow$4.78) 71.54 ($\downarrow$0.83) 72.63 ($\downarrow$2.47) 81.00 ($\downarrow$2.71)
[Uncaptioned image] Gemini-2.5-Pro w/o web search 1.32 2.29 22.56 70.51 74.48 75.52 84.51
LocalPlayground 1.89 ($\uparrow$0.57) 2.86 ($\uparrow$0.57) 26.09 ($\uparrow$3.53) 77.93 ($\uparrow$7.42) 73.50 ($\downarrow$0.98) 78.26 ($\uparrow$2.74) 83.81 ($\downarrow$0.70)
[Uncaptioned image] Qwen-Plus-Latest w/o web search 1.79 2.79 27.76 77.19 73.21 68.78 83.41
LocalPlayground 2.59 ($\uparrow$0.80) 3.12 ($\uparrow$0.33) 32.79 ($\uparrow$5.03) 80.94 ($\uparrow$3.75) 73.65 ($\uparrow$0.44) 68.68 ($\downarrow$0.10) 83.14 ($\downarrow$0.27)
[Uncaptioned image] LongCat-Large-32K w/o web search 1.59 2.49 30.93 79.45 71.82 65.58 80.72
LocalPlayground 2.73 ($\uparrow$1.14) 3.22 ($\uparrow$0.73) 33.19 ($\uparrow$2.26) 80.51 ($\uparrow$1.06) 73.40 ($\uparrow$1.58) 60.80 ($\downarrow$4.78) 81.14 ($\uparrow$0.42)
[Uncaptioned image] Hunyuan-T1 w/o web search 1.82 2.65 30.10 74.50 73.53 71.76 84.50
LocalPlayground 2.30 ($\uparrow$0.48) 3.15 ($\uparrow$0.50) 32.94 ($\uparrow$2.84) 80.59 ($\uparrow$6.09) 73.00 ($\downarrow$0.53) 63.77 ($\downarrow$7.99) 84.71 ($\uparrow$0.21)
Open-source LRMs
[Uncaptioned image] Qwen3-14B w/o web search 1.20 2.16 21.07 71.54 70.67 54.47 82.71
LocalPlayground 2.04 ($\uparrow$0.84) 2.85 ($\uparrow$0.69) 26.42 ($\uparrow$5.35) 76.25 ($\uparrow$4.71) 70.97 ($\uparrow$0.30) 48.54 ($\downarrow$5.93) 81.54 ($\downarrow$1.17)
[Uncaptioned image] Qwen3-32B w/o web search 1.33 2.10 22.74 68.23 70.40 47.73 80.80
LocalPlayground 2.48 ($\uparrow$1.15) 2.89 ($\uparrow$0.79) 24.75 ($\uparrow$2.01) 71.37 ($\uparrow$3.14) 70.17 ($\downarrow$0.23) 40.84 ($\downarrow$6.89) 81.34 ($\uparrow$0.54)
[Uncaptioned image] Qwen3-235B-A22B w/o web search 1.26 2.10 23.08 66.56 70.50 54.94 81.20
LocalPlayground 2.35 ($\uparrow$1.09) 2.81 ($\uparrow$0.71) 28.43 ($\uparrow$5.35) 73.48 ($\uparrow$6.92) 70.54 ($\uparrow$0.04) 52.43 ($\downarrow$2.51) 81.67 ($\uparrow$0.47)
[Uncaptioned image] GLM-4.5 w/o web search 1.92 2.92 29.10 76.56 67.86 73.15 80.97
LocalPlayground 2.73 ($\uparrow$0.81) 3.66 ($\uparrow$0.74) 33.78 ($\uparrow$4.68) 76.76 ($\uparrow$0.20) 71.81 ($\uparrow$3.95) 73.12 ($\downarrow$0.03) 76.66 ($\downarrow$4.31)
[Uncaptioned image] Deepseek-V3.1 w/o web search 2.36 3.36 29.77 78.64 72.17 69.25 81.56
LocalPlayground 3.43 ($\uparrow$1.07) 4.02 ($\uparrow$0.66) 34.34 ($\uparrow$4.57) 80.00 ($\uparrow$1.36) 72.39 ($\uparrow$0.22) 60.80 ($\downarrow$8.45) 81.48 ($\downarrow$0.08)
Average w/o web search 1.56 2.49 25.58 73.38 71.70 65.63 82.41
LocalPlayground 2.43 ($\uparrow$0.87) 3.13 ($\uparrow$0.64) 29.95 ($\uparrow$4.37) 77.33 ($\uparrow$3.95) 72.10 ($\uparrow$0.40) 61.99 ($\downarrow$3.64) 81.65 ($\downarrow$0.76)
Finding 2: Web search improves correctness and completeness but sacrifices faithfulness. Closed-source models outperform open-source ones in completeness and efficiency. However, overall performance remains low, highlighting the substantial challenge of multi-hop local life service search.
5.4. Sensitivity Study (RQ3)
To investigate the impact of maximum conversation rounds on tool usage and model performance, we conduct a sensitivity study using DeepSeek-V3.1, the best-performing model in correctness from RQ2, with varying max conversation round $N$ from 2 to 8.
@@ 5.4. Sensitivity Study (RQ3)
As shown in Table 6, the optimal configuration is $N=5$, which achieves the highest Correctness (34.34%). When $N<5$, insufficient rounds limit information gathering, resulting in lower correctness (30.87%–32.11%). When $N>5$, despite increased tool usage, correctness begins to decline (31.32%–32.77%), suggesting that excessive rounds introduce noisy information that hinders answer quality rather than improving it. In terms of computational cost, increasing the maximum rounds initially leads to more tool calls and longer conversations, but the growth saturates beyond $N=5$. The average tool calls plateau around 3.68–3.81 and average rounds stabilize near 4.29–4.48. Therefore, $N=5$ represents a critical turning point that balances performance and efficiency (3.43 tool calls, 4.02 rounds).
Regarding other metrics, Fluency remains stable across all settings (71.44%–72.39%), while faithfulness gradually decreases from 64.78% ($N=2$) to 60.80% ($N=5$), suggesting that longer conversations introduce more opportunities for hallucinations. Safety scores stay consistently high (80.39%–82.39%), indicating robustness across different round limits.
Table 6. Sensitivity analysis of Max Round N on model performance.
Max Round Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%)
$N=2$ 1.96 1.98 30.87 79.63 71.44 64.78 81.14
$N=3$ 2.58 2.92 31.12 79.43 71.57 63.50 82.27
$N=4$ 3.20 3.75 32.11 79.70 72.27 64.26 82.39
$N=5$ 3.43 4.02 34.34 80.00 72.39 60.80 81.48
$N=6$ 3.79 4.45 32.77 80.20 71.96 60.56 80.91
$N=7$ 3.81 4.48 31.65 80.20 71.89 61.32 80.84
$N=8$ 3.68 4.29 31.32 79.26 71.65 61.33 80.39
Finding 3: Setting maximum rounds to 5 achieves optimal performance, maximizing correctness while maintaining efficiency. Beyond this threshold, tool usage saturates but correctness declines, indicating diminishing returns from information redundancy.

### v2

@@ 5.3. Overall Performance
The choice of underlying LRM significantly influences performance across all metrics. Deepseek-V3.2 achieves the highest Correctness (35.60%), followed by GLM-4.6 (32.80%) and Gemini-2.5-Pro (32.36%). It also leads in Completeness (77.56%) and Faithfulness (39.78%). Thinking models outperform non-thinking ones in answer quality metrics (Correctness: 30.49% vs. 21.34%, Completeness: 64.59% vs. 49.10%) and most trajectory evaluation metrics , with similar Search Efficiency. Despite these variations, overall performance remains limited, with the best model achieving only 35.60% Correctness, underscoring the substantial challenge of multi-hop local search. Across all 16 models, correctness further drops on L4 tasks (23.51%) compared to L3 (28.66%), reflecting the greater challenge of L4 (see Appendix A.6).
Human evaluation. While LLM-as-judge provides scalable assessment, it may introduce biases. Our human evaluation validates its reliability, showing high consistency with human judgment in relative model ranking. We conduct two studies: (1) 3 annotators evaluate 900 outputs from Deepseek-V3.2, and (2) 100 randomly sampled outputs each from GLM-4.6, Gemini-2.5-Pro, LongCat-Flash-Thinking, and Qwen3-235B-A22B. Human scores demonstrate substantial inter-rater reliability (Cohen’s $\kappa$ = 0.79, 95% CI: $[0.75,0.83]$, 87.91% raw agreement). Detailed results are in Appendix A.2.
5.4. Ablation Study
We conduct ablation study on tool integration–a core design element for agentic search in local life services. Specifically, we evaluate three configurations across top-performing LRMs (Gemini-2.5-Pro, GLM-4.6, Deepseek-V3.2): (1) LRM-only (no tools), (2) LRM + LocalRAG (merchant retrieval only), and (3) LRM + LocalRAG + web search (full toolchain). The goal is to quantify how each tool impacts answer and trajectory quality.
As shown in Table 7, incorporating web search into LocalPlayground improves answer quality but reduces faithfulness. On average, it boosts Correctness by 8.90 pp and Completeness by 7.41 pp, while decreasing Faithfulness by 2.83 pp. This trade-off is most evident in Gemini-2.5-Pro, which gains 9.25 pp in Correctness but loses 3.72 pp in Faithfulness. These results highlight web search’s dual role in enabling critical information retrieval and introducing noise in multi-hop reasoning.
5.5. Sensitivity Study
@@
To investigate the impact of maximum conversation rounds $N$ on tool usage and model performance, we conduct a sensitivity study using Deepseek-V3.2, the best-performing model, with varying $N$ from 2 to 8. As shown in Table 8, the optimal configuration is $N=5$, achieving the highest Correctness (35.60%). When $N<5$, insufficient rounds limit information gathering, resulting in lower correctness (30.50%–34.20%). When $N>5$, despite increased tool usage, correctness declines (34.90%–32.48%), suggesting excessive rounds introduce noise that hinders answer quality. In terms of computational cost, tool calls and conversation length increase with $N$, peaking at $N=8$ (4.02 tool calls, 5.01 rounds). Therefore, $N=5$ represents the optimal balance of performance and efficiency.
5.6. In-Depth Analysis
Figure 9. Failure Case Studies. A typical chain event: the 2nd and 3rd hops depend on the 1st hop.
We study factors affecting LRM performance on agentic search tasks in local life services. Figure 9 shows concrete examples.
Unstable tool call strategy (38.6%). Lack of a unified strategy for tool call timing and frequency leads to insufficient tool calls or wasted budget. LRMs may also violate tool-use rules, such as using web search for merchant queries or making redundant tool calls within a single dialogue turn.
Missing explicit multi-hop planning (30.9%). LRMs lacking hop sequences fail to decompose tasks for retrieval or tool use. Without explicit hop targets, sub-goals remain unordered, causing empty matches and higher coverage failures in multi-hop tasks. Retrieval failures cascade through steps due to the absence of validation or backtracking.
Query generation dilutes critical constraints (11.8%). Naively generated long-form queries often omit key constraints, weakening retrieval signals and reducing recall of relevant candidates.
Long-context noise (18.7%). Lengthy search results and information sent to LRMs bring irrelevant data, reduce the ability to distinguish relevant results, and can exceed context window limits.